# Group Testing with General Correlation Using Hypergraphs

### v1

@@ 9 Else:
Let’s run Algorithm 1 on the graph in Figure 1 with $c=0.1$. Initially, $S=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$ and $w(S)=1$. By removing $v_{1}$ from $S$, $E(S\setminus v_{1})$ contains only one edge $e_{3}=\{4,5\}$, hence $c<w(S\setminus v_{1})=0.5<1-c$ and the algorithm tests $v_{1}$. If it is negative, then it rules out $e_{1}$ and $e_{2}$ and after scaling the edges, $q_{e_{3}}=1$ and the algorithm returns it. If it is positive, it removes $e_{3}$ and scales the other edges by 2 so $q_{e_{1}}=0.6$ and $q_{e_{2}}=0.4$. Again $S$ would be initialized to contain all the nodes, and after removing $v_{2}$, $w(S\setminus v_{2})=0.4$ so the algorithm test $v_{2}$ to eliminate one of $e_{1}$ or $e_{2}$. The expected number of tests here is 1.5.
In the following, we give a more abstract and larger example.
@@
To give another illustrating example, suppose there are $k$ islands where each island has $m$ nodes, so $n=mk$. Nodes on the same island are perfectly correlated, ie, they always have the same state, as such every hyper edge always contains all or none of any given island (Figure 4, $k=4,m=3$). Thus, there are $2^{k}$ hyperedges at most; note that in this case the number of hyper edges can be significantly smaller than than the size of the power set of nodes, $2^{mk}$. Consider a setting with three types of islands: low, moderate and high, characterized by their probability of infection. For all nodes in high islands, $p_{i}\approx 1$ (e.g. island 4 in Figure 4). In moderate islands (e.g. islands 2,3), $p_{i}\approx\frac{1}{2}$ and in low islands, e.g. island 1, $p_{i}\approx\epsilon$, where $\epsilon$ is a small, positive number. One can verify that Algorithm 1 first picks a candidate from one of the moderate islands in line 1 and tests it individually. If it is positive, posterior probabilities of all nodes in its island is updated to $1$, otherwise to 00, through an update of posterior probabilities of all the edges (line 1 or 1). Accordingly, no other node from this island is identified for testing in line 1. Subsequently, similarly one node from each moderate island is identified for testing in line 1, and posterior probabilities are updated which eliminates the need to test other nodes from the same island in line 1. At the end of this process, all islands are high or low. Then, the algorithm gathers all the nodes from all low islands, in line 1 and tests them at once in line 1. With high probability, the test is negative, and finally, the algorithm tests nodes in high islands individually, except those whose posterior probability is already updated to $1$, updates the posterior probabilities of other nodes after each test, and returns all the nodes with $q_{v}=1$. Note that algorithm operates such that only one node is tested from each island.
Figure 4: An example with $k=4$ islands. Each edge contains either all nodes in an island or none of it. Three edges are given in red. Nodes in island 1 are infected with $p_{1}\approx 0$, nodes in islands 2 and 3 are infected with probability $p_{2}=p_{3}\approx 1/2$, and nodes in island 4 are infected with high probability $p_{4}\approx 1$.
@@ 4 Analysis
In this section, we bound the expected number of tests performed by Algorithm 1. The main result is built on the following key lemma. Here, $H(X)$ is the entropy of the nodes and $\tilde{\mu}$ is the expected number of infections in line 1 (computed with the posterior probabilities).
@@
Algorithm 1 finds $e^{*}$ with $\mathbb{E}[L_{1}+L_{2}]\leq\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{\mathbb{E}[{\tilde{\mu}}]}{1-2c}$ tests in expectation.
@@
Informative Tests in Stage 1.
The above lemma implies that if both $E(S)$ and $E\setminus E(S)$ have a moderate mass (i.e., $w(S)\geq c$ and $1-w(S)\geq c$), regardless of the result, the test is informative. This is quantified next for Stage 1 of the algorithm. We first give a structure on probabilities of nodes in line 1, where the algorithm fails to find a set $S$ so that $c\leq w(S)\leq 1-c$.
Proposition 4.4.
When Algorithm 1 reaches  1, we have (i) $w(S)>1-c$, (ii) $Pr(\exists v\in V\setminus S:v\in e^{*})<c$ where $e^{*}$ is the sampled edge, and (iii) $\forall v\in S:q_{v}>1-2c$.
@@
To see (i) note that at line 1 $w(S)=1>1-c$ and, as we proceed by removing nodes from $S$ to meet the criteria of Line 1, it never gets met and therefore when the algorithm reaches Line 1, we have $w(S)>1-c$. To prove (ii), we compute the probability that at least one node $v\in V\setminus S$ is positive. By Lemma 4.2, this is equal to the probability that $e^{*}\in E\setminus E(S)$, i.e., $\sum_{e\in E\setminus E(S)}q_{e}=1-w(S)$ which is at most $c$ by (i). Finally, to prove (iii), we note that $v\in S$, $q_{v}\geq w(S)-w(S\setminus v)$, because some of edges that contain $v$ have nodes outside of $S$. But the algorithm has passed line 1, so $w(S)>1-c$ and $w(S\setminus v)<c$ for any node $v$. Therefore, $q_{v}\geq w(S)-w(S\setminus v)>1-2c$.
The following observation shows that the tests in the first stage are “informative”.
@@
For Algorithm 1, in Stage 1, for every update of $\mathcal{D}$ (lines 1, 1,1), at least $c$ fraction of the mass is removed.
@@
If the algorithm scale $\mathcal{D}$ at lines 1 or 1, by lemma 4.2, we either remove $E(S)$ or $E\setminus E(S)$. But $c\leq w(S)\leq 1-c$, so in either case we remove $c$ fraction of the mass.
If the algorithm scales $\mathcal{D}$ at line 1, it means the test had been positive. By Proposition 4.4, $w(S)=1-Pr(v\in V\setminus S\cap v\in e^{*})>1-c>c$ because $1-c>1/2>c$, meaning at least $c$ mass have been removed. So for any update, at least $c$ fraction is removed. ∎
Now we can bind the number of tests $L_{1}$ performed in the first stage of the algorithm using the fact that each test before line 1 is informative.
@@
$\mathbb{E}[L_{1}]\leq\frac{1}{\log\ \frac{1}{(1-c)}}H(X).$
@@
By Observation 4.5, we know that every time the posterior $\mathcal{D}$ is updated in Stage 1, at least a mass $c$ is removed and the remaining edge probabilities are scaled by at least a factor of $\frac{1}{1-c}$. Suppose that the target edge $e^{*}$ is selected with probability $p_{e^{*}}$. It takes at most $\log_{\frac{1}{1-c}}(1/p_{e^{*}})$ tests for the posterior on $e^{*}$ to become 1. So the number of tests before line 1 is at most $\log_{\frac{1}{1-c}}(1/p_{e^{*}})=\frac{\log 1/p_{e^{*}}}{\log(\frac{1}{1-c})}$, leading to the following bound on $\mathbb{E}[L_{1}]$:
$\mathbb{E}[L_{1}]<\sum_{e\in E}p(e)\log_{\frac{1}{1-c}}(p(e))\\ =\sum_{e\in E}p(e)\frac{\log 1/p(e)}{\log\frac{1}{1-c}}\\ =\frac{1}{\log\frac{1}{1-c}}\sum_{e\in E}p(e)\log 1/p(e)\\ =\frac{1}{\log\ \frac{1}{(1-c)}}H(X).$
Individual Testing in Stage 2.
Algorithm 1 stops performing in Stage 1 when the greedy process of finding $S$, $c\leq w(S)\leq 1-c$, fails in line 1, and the nodes out of $S$ tests negative in line 1. At this point, in Stage 2, we resort to individual testing in Line 1.
Using Proposition 4.4, we bind the expectation of $L_{2}$ based on the expectation at the “stopping time” $\tilde{\mu}$ in the following.
@@
Let $\tilde{\mu}=\mu_{(\mathbf{T_{1},r_{1}),\ldots,(T_{L_{1}},r_{L_{1}})}}$ be the expected number of infected nodes at the end of line 1 where $\mathbf{(T_{i},r_{i})},1\leq i\leq L_{1}$ are random variables which indicate the tests and results performed. Then the algorithm performs at most $L_{2}\leq\frac{\tilde{\mu}}{1-2c}$ tests in line 1.
@@
Recall that $S$ here is defined in Algorithm 1 when it reached Line 1. The first equality is true because in the previous line, $V\setminus S$ tested negative, so $X_{v}=0$ for $v\in V\setminus S.$ The second inequality is true because given the test results and which nodes have been tested thus far, the set $S$ is specified, thus the expectation can be moved inside the summation. The third equality follows from what $q_{v}$ stands for. The fourth inequality follows from Proposition 4.4, from which we know that for each $v\in S$, $q_{v}>1-2c$. But the algorithm performs $|S|$ tests at line 1, so $L_{2}=|S|$. Thus regardless of the test results, $L_{2}<\frac{\tilde{\mu}}{1-2c}.$
The above observation in conjunction with Corollary 4.6 proves Lemma 4.1.
@@
If $\forall e\in E:f_{1}(n)\leq|e|\leq f_{2}(n)$, then Algorithm 1 finds $e^{*}$ with $\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{f_{2}(n)}{f_{1}(n)(1-2c)}\mu$ tests in expectation.
@@
Recall that $\mu$ is the average number of infections before performing any tests. As $f_{1}(n)\leq|e|$ for each edge $e$, we have $\mu\geq f_{1}(n)$. As $|e|\leq f_{2}(n)$, we have $\tilde{\mu}\leq f_{2}(n)$. Then the ratio $\frac{\tilde{\mu}}{\mu}\leq\frac{f_{2}(n)}{f_{1}(n)}$, hence $\mathbb{E}[\tilde{\mu}]\leq\frac{f_{2}(n)}{f_{1}(n)}\mu$. Replacing this in $\mathbb{E}[L_{1}+L_{2}]=\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{\mathbb{E}[\tilde{\mu}]}{1-2c}$ from Lemma 4.1 completes the proof. ∎
Note that if most of the edges have size concentrated around $\mu$, we can remove very large or very small edges and incur a small probability of error, but get a better bound. This idea is summarized in the following corollary.
@@
If there is an $E^{\prime}\subseteq E$ such that $\sum_{e\in E^{\prime}}p(e)=1-\epsilon$ and $\forall e\in E^{\prime}:f_{1}(n)\leq|e|\leq f_{2}(n)$, then Algorithm 1 finds $e^{*}$ with $\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{f_{2}(n)}{f_{1}(n)(1-2c)}\mu$ tests in expectation with probability $1-\epsilon$.
While Theorem 4.8 imposes a constraint on the edge sizes and finds the infected set with probability 1, the constraint can be relaxed if a small probability of error ($\epsilon$, fixed or approaching zero) can be tolerated. We use Markov Lemma to bound the probability that $|e^{*}|$ exceeds $\mu/\epsilon$. By removing those edges for which $|e|>\mu/\epsilon$, we ensure $\tilde{\mu}<\mu/\epsilon$ and establish the following result.
@@
There is an algorithm that performs $\mathbb{E}[L_{1}+L_{2}]\leq\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{\mu}{\epsilon(1-2c)}$ tests in expectation and returns the correct edge $e^{*}$ with probability $1-\epsilon$.

### v2

@@ 10 Else:
We now run Algorithm 1 on the graph in Figure 1 with $c=0.1$. Refer to the values of $\{q_{e}\}$ in the caption of Figure 1, $\{p_{v}\}$ calculated in Section 3.2 and that $w(V\setminus\{v\})=1-p_{v}$ also from Section 3.2. In line 1 we set $S=V=\{v_{1},v_{2},v_{3},v_{4},v_{5}\}$, where $W=\phi$. Thus, $w(S)=1$. Thus, $S$ does not satisfy the condition in line 1, and the Algorithm moves to line 1 and then line 1. Here, $w(V\setminus\{v_{1}\})=1-p_{v_{1}}=0.5$. Thus, $c<w(S\setminus v_{1})=0.5<1-c$, the algorithm removes $v_{1}$ from $S$, and tests $v_{1}$ in line 1. If it is negative, then $v_{1}$ is added to $W$, $e_{1}$ and $e_{2}$ are ruled out as target edges as both contain $v_{1}$, and after scaling the edges in line 1 following Definition 3.2, posterior $q_{e_{3}}=1$ since $e_{3}$ is the only remaining edge. And the algorithm returns it as the target edge in line 1. If $v_{1}$ is positive, in line 1, $e_{3}$ is ruled out as the target edge as it does not contain $v_{1}.$ Since $q_{e_{3}}=0.5$, in line 1, the algorithm scales $q_{e_{1}},q_{e_{2}}$ by 2 so posterior $q_{e_{1}}=0.6$ and $q_{e_{2}}=0.4$. Again in line 1, $S$ would be initialized to $V.$ Thus $w(S)=1$ and $S$ does not satisfy the criterion in line 1. Thus, the algorithm moves to line 1, finds that there is no $q_{e}$ such that $q_{e}=1.$ It then moves to line 1. Since $S=V$, $w(S\setminus\{v_{2}\})=1-p_{v_{2}}=0.4$, so the algorithm identifies $v_{2}$ for individual test. The test result will eliminate $e_{1}$ if $v_{2}$ is positive, and $e_{2}$ otherwise in lines 1 and 1. Either way, only one edge remains, thus its posterior probability of being the target edge is scaled to $1$ in line 1 or 1 (Definition 3.2), and it is identified as the target edge in line 1. The expected number of tests here is $1.5$
In the following, we give an important remark and proceed with a more abstract and larger example.
Remark 3.5.
We remark that any node for which the posterior probability becomes 0 or 1 will no longer be tested. In particular, if an individual node (set) is tested negative, since it is removed from $V$, it is never tested again. If an individual node (set) is tested positive, only edges intersecting that node (set) remain in the graph, thus the weight of the complement of that node (set) becomes 00. Thus, the complement set will never have weight in the interval $[c,1-c]$ and is therefore not tested again. In the following, we elaborate in details on why a single node is not tested again once its infection probability reaches zero or one.
Suppose at any stage of Algorithm 1 all remaining edges contain a particular node $v.$ This may happen for example right after $v$ tests positive in an individual test. Then following Definition 3.2, the posterior probability of the node will be $1$, and removing the node from any set $S$ will result in $S$ not containing any edge, and the weight of the set will become 00. Thus, such a node can not be included in any test.
In addition, if at any stage of the algorithm no remaining edge contains a node $v$, which can for example happen if any group including $v$ tests negative, then its posterior probability becomes 00, and it is moved to $W$ from $V.$ Right after a node is moved to $W$, the set $S$ is initialized as $V\setminus W$ in line 1, thus such a node does not belong to $S$ henceforth and is therefore not removed from it in any line and is therefore not tested further.
@@
We now elucidate how Algorithm 1 functions for Example 3.3 in Section 3.2. First note that whenever a node $v$ is tested individually, the test reveals its state and also the states of all nodes in its island as all nodes in an island have the same state. Accordingly no node in $v$’s island is tested further as we argue next. If $v$ is negative, all edges that contain it are removed. But any edge containing it also contains all nodes in its island. Thus all edges containing any node in $v$’s island are removed. If $v$ tests positive, only the edges containing it are retained (line 1), and all such edges contain all the nodes in its island since every edge contains all or none of the nodes in each island. Thus, every remaining edge has all nodes in $v$’s island. Thus, by Remark 3.5, no node in $v$’s island is tested henceforth regardless of the outcome of $v$’s test. We now describe the sequence of testing of nodes.
The first decision on which node to test is made when Algorithm 1 reaches line 1 for the first time. Then $S=V$ and $w(S)=1.$ Recall that we had argued in Section 3.2 that $w(V\setminus\{v\})=1-p_{v}.$ Thus, $w(S\setminus\{v\})$ is $\gamma$, $\approx 1/2,$ $1-\epsilon$ if $v$ is respectively in a high, moderate, low island, where $\gamma,\epsilon$ are small positive numbers.
Assuming that $\gamma,\epsilon<c$, it follows that when the algorithm reaches line 1 for the first time, $c<w(S\setminus v)<1-c$ if and only if $v$ is in a moderate island. Thus, Algorithm 1 first picks a candidate from one of the moderate islands in line 1 and tests it individually. As discussed above, testing $v$ reveals the state of all nodes in $v$’s island and subsequently no other node from the same island is tested. Since nodes in different islands are independent, the knowledge of the state of the tested node does not affect the probabilities that nodes in other islands are positive. Thus following the same arguments as before the algorithm continues to visit line 1, and in each such visit identifies a node from a new moderate island for individual testing. The outcome of the test reveals the state of all nodes in its island and subsequently no other node from the same island is tested. This continues until all moderate islands are exhausted. The remaining islands are high or low.
Then, the algorithm reaches line 1 and identifies a low probability node there, that is, a node from a low island. Assuming that the $\epsilon<c/k_{3}$ where $k_{3}$ is the number of low islands, removing an additional node in line 1, either reduces the weight of $S$ to below $c$ or maintains it at greater than $1-c$, thus no further node is identified for testing in line 1.
Repeated visits to line 1, gathers all the nodes from all low islands, and tests them all together in line 1. With high probability, the test is negative, and if so, all nodes in the low islands are moved to $W$, all edges containing low islands are removed, and only the high islands remain. Then the algorithm tests nodes in high islands individually in line 1. Each individual test reveals the state of all nodes in the tested node’s island and no other node in the tested node’s island need to be tested further as argued in the first paragraph. The algorithm outputs the edge containing all the positive nodes as the target edge.
If however the group test of all the nodes in the low islands turn out to be positive, which happens with low probability, then these nodes would subsequently be either tested individually, identified one at a time in line 1, or tested individually in line 1, or tested in smaller subgroups which are identified in line 1 or in both line 1 and line 1.  [^1] Here, we describe in details how the algorithm performs in the unlikely event that low islands test positive. If the group test of all the nodes in the low islands turns out to be positive, the edges consisting only of the high islands will be removed. This would substantially increase the posterior probabilities of edges which include low islands which would in turn substantially increase the posterior probabilities of the nodes in low islands. The posterior probabilities of the nodes in high islands will not change because of independence of states of nodes across islands. The algorithm will move to line 1 and subsequently to line 1. Since the posterior probabilities of nodes in high islands remain same as before, they will not qualify for individual tests as before. But, since the posterior probabilities of nodes in low islands have increased, they may qualify for individual tests in line 1. If so, as for moderate islands, one node in each low island would be tested and the outcome would be used to determine the states of each node in the same island and those nodes will not be tested further. At the end of the process, only the high islands will remain. Since they are all high probability nodes, no node would be identified for individual testing in line 1. But these nodes would be individually tested in line 1 and their states determined. If however the posterior probabilities of nodes in low islands increased significantly such that the removal of any of them reduces the weight of $S$ to below $c$, then they would not qualify for individual tests in line 1. Then all remaining nodes including those of low islands are high probability nodes. Thus they would not be removed in line 1 either, and all remaining nodes would be individually tested in line 1. The last possibility is that the posterior probabilities of nodes in low islands increased, but still removing any one of them retains the weight of $S$ to above $1-c.$ Then, again they would not qualify for individual tests in line 1. But they can be removed as low probability nodes in line 1. But note that all nodes of low islands can not be removed together as low probability nodes in line 1, because only edges that contain at least one of them have been retained, thus together all remaining edges contain all of them, and hence removing all of them removes all the edges too and reduces the weight of $S$ to $0<c.$ Then subsets of nodes in low islands would be identified in lines 1 and 1 and tested in line 1, and the sequence is repeated.
Summarily, the nodes in high and moderate islands are individually tested, but those in low islands are tested all together at least once, and in the unlikely scenario that this group test is positive they are tested in various subgroups to determine their states. Only one node in each high and moderate island is tested.
Remark 3.7.
In line 1, the algorithm identifies the first node that satisfies the condition given in this line, and removes it from the set $S$. However, the algorithm could instead consider all nodes individually or multiple nodes simultaneously. Doing so would ensure not only that the condition in this line is satisfied but also that the weight of the set $S$ after removal is as close to $1/2$ as possible. A set weight closer to $1/2$ results in a more informative test in practice. However, such a selection strategy, while potentially improving practical performance, would be computationally more expensive and would not affect the theoretical guarantees or our analysis.
@@ 4 Analysis
In this section, we bound the expected number of tests performed by Algorithm 1. The main result is built on the following key lemma. Here, $H(X)=\sum_{e}p_{e}\log 1/p_{e}$ is the entropy of the nodes and $\tilde{\mu}$ is the expected number of infections in line 1 (computed with the posterior probabilities).
@@
Algorithm 1 finds $e^{*}$ with $\mathbb{E}[L_{1}+L_{2}]\leq\frac{1}{\log\frac{1}{(1-c)}}H(X)+1+\frac{\mathbb{E}[{\tilde{\mu}}]}{1-2c}$ tests in expectation.
@@
Informative Tests:
Formally, a test is informative if the edge updates right after it remove at least $c$ fraction of the mass. The following lemma upper bounds the expected number of informative tests for any algorithm that returns an edge with weight $1$ as the target edge.
Lemma 4.4.
The expected number of informative tests is at most $\frac{1}{\log\ \frac{1}{(1-c)}}H(X)$ for any algorithm that decides that an edge $e$ is the target edge whenever $q_{e}=1.$
@@
We know that every time the posterior is updated after every informative test, at least a mass $c$ is removed and the remaining edge probabilities are divided by at most a factor of $1-c$. It therefore takes at most $\log_{\frac{1}{1-c}}(1/p(e))=\frac{\log 1/p(e)}{\log(\frac{1}{1-c})}$ informative tests for the posterior on a given edge $e$ to become 1 (if it survives those tests which it would if it is the target edge). This leads to the following upper bound on the expected number of informative tests:
$\sum_{e\in E}p(e)\log_{\frac{1}{1-c}}(p(e))\\ =\sum_{e\in E}p(e)\frac{\log 1/p(e)}{\log\frac{1}{1-c}}\\ =\frac{1}{\log\frac{1}{1-c}}\sum_{e\in E}p(e)\log 1/p(e)\\ =\frac{1}{\log\ \frac{1}{(1-c)}}H(X).$
The following observation identifies the informative tests in the algorithm.
@@
For Algorithm 1, tests in line 1 and positive tests in line 1 are informative, provided $20<c<1/$ 2 .
@@
It would be sufficient to prove that edge updates in lines 1, 1,1 remove at least $c$ fraction of the mass. If the algorithm scale $\mathcal{D}$ at lines 1 or 1, by lemma 4.2, we either remove $E(S)$ or $E\setminus E(S)$. But $c\leq w(S)\leq 1-c$, so in either case we remove $c$ fraction of the mass.
If the algorithm scales $\mathcal{D}$ at line 1, it means $w(S)>1-c$ and $V\setminus S$ has tested positive. The latter implies that $E(S)$ is removed from $E.$ Weight of $E(S)$ is $w(S).$ Thus, at least a mass of $1-c$ is removed. Now $1-c>c$ because $1-c>1/2>c$. Thus at least a mass of $c$ have been removed. ∎
Note that there can be at most one negative test in line 1, since right after such a negative test, the algorithm proceeds to line 1, and never returns to lines $1$ to 1. Thus, all but $1$ tests performed before the algorithm proceeds to line 1, ie, in Stage $1$ of the algorithm, are informative. Using Lemma 4.4, now we can upper bound the expected number of tests performed in the first stage of the algorithm.
@@
$\mathbb{E}[L_{1}]\leq\frac{1}{\log\ \frac{1}{(1-c)}}H(X)+1.$
Line 1 constitutes Stage $2$ of the algorithm.
Individual Testing in Stage 2.
Stage $1$ of Algorithm 1 concludes when the greedy process of finding $S$, $c\leq w(S)\leq 1-c$, fails in line 1, and the nodes outside $S$ test negative in line 1. At this point, in Stage 2, we resort to individual testing in line 1. We will upper bound the expected number of such individual tests using the following Proposition.
Proposition 4.7.
When Algorithm 1 reaches 1, we have (i) $w(S)>1-c$, (ii) $Pr(\exists v\in V\setminus S:v\in e^{*})<c$ where $e^{*}$ is the target edge, (iii) $\sum_{e\in E(S),v\in e}p_{e}>1-2c$, and (iv) $\forall v\in S:q_{v}>1-2c$.
@@
To see (i) note that at line 1 $w(S)=1>1-c$ and, as we proceed by removing nodes from $S$ to meet the criteria of line 1, it never gets met and therefore when the algorithm reaches line 1, we have $w(S)>1-c$. To prove (ii), we compute the probability that at least one node $v\in V\setminus S$ is positive. By Lemma 4.2, this is equal to the probability that $e^{*}\in E\setminus E(S)$, i.e., $\sum_{e\in E\setminus E(S)}q_{e}=1-w(S)$ which is at most $c$ by (i). Finally, to prove (iii) and (iv), we note that $\forall v\in S$, $q_{v}\geq\sum_{e\in E(S),v\in e}q_{e}=w(S)-w(S\setminus v)$. Since the algorithm has passed line 1, so $w(S)>1-c$ and $w(S\setminus v)<c$ for any node $v$. Therefore, $q_{v}\geq\sum_{e\in E(S),v\in e}q_{e}>1-2c$, proving (iii) and (iv).
Using Proposition 4.7, we bound the expected number of tests in line 1:
@@
Let $\tilde{\mu}=\mu_{(\mathbf{T_{1},r_{1}),\ldots,(T_{L_{1}},r_{L_{1}})}}$ be the expected number of infected nodes at the start of line 1 where $\mathbf{(T_{i},r_{i})},1\leq i\leq L_{1}$ are random variables which indicate the tests and results performed. Then the algorithm performs at most $L_{2}\leq\frac{\tilde{\mu}}{1-2c}$ tests in line 1.
@@
Recall that $S$ here is defined in Algorithm 1 when it reached line 1. The first equality is true because in the previous line, $V\setminus S$ tested negative, so $X_{v}=0$ for $v\in V\setminus S.$ The second equality is true because given the test results and which nodes have been tested thus far, the set $S$ is specified, thus the expectation can be moved inside the summation. The third equality follows from what $q_{v}$ stands for. The fourth inequality follows from Proposition 4.7 (iii), from which we know that for each $v\in S$, $\sum_{e:e\in E(S),v\in e}p_{e}>1-2c$ in line 1. As the test result in line 1 is negative, following Definition 3.2, the weight of the edges in $E(S)$ only scaled up at the start of line 1. Thus, $\sum_{e:e\in E(S),v\in e}p_{e}>1-2c$, right after the negative test. Since $\forall v:q_{v}\geq\sum_{e:e\in E(S),v\in e}p_{e}$, $q_{v}>1-2c$. But the algorithm performs $|S|$ tests at line 1, so $L_{2}=|S|$. Thus regardless of the test results, $L_{2}<\frac{\tilde{\mu}}{1-2c}.$
The above lemma in conjunction with Corollary 4.6 proves Lemma 4.1.
@@
If $\forall e\in E:f_{1}(n)\leq|e|\leq f_{2}(n)$, then Algorithm 1 finds $e^{*}$ with $\frac{1}{\log\frac{1}{(1-c)}}H(X)+\frac{f_{2}(n)}{f_{1}(n)(1-2c)}\mu+1$ tests in expectation.
@@
Recall that $\mu$ is the average number of infections before performing any tests. As $f_{1}(n)\leq|e|$ for each edge $e$, we have $\mu\geq f_{1}(n)$. As $|e|\leq f_{2}(n)$, we have $\tilde{\mu}\leq f_{2}(n)$. Then the ratio $\frac{\tilde{\mu}}{\mu}\leq\frac{f_{2}(n)}{f_{1}(n)}$, hence $\mathbb{E}[\tilde{\mu}]\leq\frac{f_{2}(n)}{f_{1}(n)}\mu$. Replacing this in $\mathbb{E}[L_{1}+L_{2}]\leq\frac{1}{\log\frac{1}{(1-c)}}H(X)+1+\frac{\mathbb{E}[\tilde{\mu}]}{1-2c}$ from Lemma 4.1 completes the proof. ∎
If most of the edges, not all edges, have an upper bound of the form $f_{2}(n)$, Algorithm 1 can be modified to only return the correct edge when the target edge is smaller or equal than $f_{2}(n)$. This idea is summarized in the following corollary.
@@
If there is an $E^{\prime}\subseteq E$ such that $\sum_{e\in E^{\prime}}p(e)=1-\epsilon$ and $\forall e\in E^{\prime}:|e|\leq f_{2}(n)$, then Algorithm 1 can be modified so that it performs $\frac{2}{\log\frac{1}{(1-c)}}H(X)+2f_{2}(n)$ tests in expectation, and returns the target edge $e^{*}$ with probability $1-\epsilon$, where $c<1/3$.
Proof.
Let $PN$ be the number of positive tests in line 1, which is modified as follows:
  1. 16.
    If the test is negative, remove $E\setminus E(S)$ from $\mathcal{D}$ and update the posterior distribution following Definition 3.2. Pick a node $v\in S$ at random with $10<q_{v}<$ 1. Test $v$ and update the posterior and if $v$ is positive increment $PN$ by $1$. If $PN\geq f_{2}(n)$, return all the positive nodes as $e^{*}$; else, go to line 1.
First consider the informative tests. As per Lemma 4.4, the expected number of informative tests is at most $\frac{1}{\log\ \frac{1}{(1-c)}}H(X)$. As in Observation 4.5, tests in line 1 and positive tests in line 1 are informative. We now prove that a negative test in line 16 is informative. When the algorithm reaches line 1 , for each $v$ in $S$, $\sum_{e:e\in E(S),v\in e}p_{e}>1-2c$ (this can be argued similar to Proposition 4.7). If a node $v$ tests negative in line 16, then all edges that contain $v$ are removed, thus a weight of at least $\sum_{e:e\in E(S),v\in e}p_{e}$ is removed (note that edges that are not contained in $E(S)$ but have $v$ will also be removed, so we use the qualifier “at least”). Thus, a weight of at least $1-2c$ is removed. Note that $1-2c>c$ since $c<1/3$. The claim follows. Thus, the expected total number of tests in line 1 and positive tests in line 1 and negative tests in line 16 is $\frac{1}{\log\ \frac{1}{(1-c)}}H(X)$.
We now upper bound the remaining tests in expectation. The remaining tests are positive tests in line 16 and negative tests in line 1. We first consider the positive tests in line 16. Here, every positive test increments $PN$ by at least $1$ It therefore follows from line 16 that there are at most $f_{2}(n)$ such positive tests. Next, note that between any two negative tests in line 1, at least one test is conducted in line 16, which is informative if negative, or positive otherwise. Thus the number of negative tests in line 1 is at most the number of informative tests and the number of positive tests in line 16. Thus, the expected number of negative tests in line 1 is at most $\frac{1}{\log\ \frac{1}{(1-c)}}H(X)+f_{2}(n).$ Putting all together, the algorithm performs at most $\frac{2}{\log\frac{1}{(1-c)}}H(X)+2f_{2}(n)$ tests in expectation.
If $|e^{*}|\leq f_{2}(n)$, we always have $PN\leq f_{2}(n)$ and hence the algorithm terminates with $PN<f_{2}(n)$ or $PN=f_{2}(n).$ In the former case the algorithm terminates only when it discovers an edge $e^{*}$ such that $p(e^{*})=1.$ In the latter case it terminates returning $f_{2}(n)$ nodes that tested positive. Since $|e^{*}|\leq f_{2}(n)$ and all positive nodes are in $e^{*}$, the algorithm returns all the positive nodes, ie, it returns $e^{*}$. But the target edge is less than or equal to $f_{2}(n)$ with probability $1-\epsilon$; hence the probability of returning the target edge is at least $1-\epsilon$ (the target edge may be returned correctly even when it exceeds $f_{2}(n)$).
Note that under the conditions of Corollary 4.10, and when $\forall e\in E^{\prime}:f_{1}(n)\leq|e|\leq f_{2}(n)$, $\epsilon<2c$, from Corollary 4.10, the expected number of tests in the modified algorithm is also at most $\frac{2}{\log\frac{1}{(1-c)}}H(X)+\frac{2f_{2}(n)}{f_{1}(n)(1-2c)}\mu$ in expectation. This is because $\mu\geq f_{1}(n)(1-2c)$ which we argue next. Note that $\mu=\mathbb{E}[|e^{*}|]$, and $|e^{*}|\geq f_{1}(n)$ with probability $1-\epsilon$. Thus, $\mu=\mathbb{E}[|e^{*}|]\geq f_{1}(n)(1-\epsilon)\geq f_{1}(n)(1-2c)$ since $\epsilon<2c$. The claim follows.
While Theorem 4.9 imposes a constraint on the edge sizes and finds the infected set with probability 1, the constraint can be relaxed if a small probability of error ($\epsilon$, fixed or approaching zero) can be tolerated. We use Markov Inequality to bound the probability that $|e^{*}|$ exceeds $\mu/\epsilon$. By removing those edges for which $|e|>\mu/\epsilon$, we ensure $\tilde{\mu}<\mu/\epsilon$ and establish the following result.
@@
There is an algorithm that performs $\mathbb{E}[L_{1}+L_{2}]\leq\frac{2}{\log\frac{1}{(1-c)}}H(X)+\frac{2\mu}{\epsilon}$ tests in expectation and returns the correct edge $e^{*}$ with probability $1-\epsilon$, where $c\leq 1/3$. Optimizing over $c$, with $c=1/3$ the expected number of tests is $\mathbb{E}[L_{1}+L_{2}]\leq\frac{2}{\log\frac{3}{2}}H(X)+\frac{2\mu}{\epsilon}$.